OpenStudy (anonymous):

divide 20x²y^5 -------- 12xy^7

OpenStudy (rulnick):

20/12 = 5/3 x^2 / x = x y^5 / y^7 = y^-2 so answer is 5 x y^-2 / 3.

OpenStudy (rulnick):

Or, if you prefer, 5x / (3 y^2)